OpenStudy (anonymous):

Jessica used Fraction 2 over 3 yard of fabric to make a scarf. Can she make 2 of these scarves with Fraction 1 and 3 over 4 yards of fabric, and why?

OpenStudy (anonymous):

Yes, because the quotient of Fraction 1 and 3 over 4 ÷ fraction 2 over 3 is 2 and 5 over 8 No, because the quotient of Fraction 2 over 3 ÷ Fraction 1 and 3 over 4 is 1 and 2 over 12 No, because the quotient of Fraction 1 and 3 over 4 ÷ fraction 2 over 3 is 1 and 2 over 12 Yes, because the quotient of Fraction 2 over 3 ÷ Fraction 1 and 3 over 4 is 2 and 5 over 8

OpenStudy (anonymous):

Yes\[1\frac{ 3 }{ 4 }\div \frac{ 2 }{ 3}\]Change the mixed number to a fraction and divide by 2/3

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }\div \frac{ 2 }{ 3}\] Invert the divisor and multiply

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }*\frac{ 3 }{ 2 }=\frac{ 21 }{ 8 }\]

OpenStudy (anonymous):

Reduce the fraction and change to mixed number.\[\frac{ 21 }{ 8 }=2\frac{ 5 }{ 8}\]